OpenStudy (anonymous):

Not to be "that" guy, but I have legitimately no clue how to do these problems... Write the equation of the line with the given slope,-2 and y-intercept, 5. A. y = -5x + 2 B. y = -2x + 5 C. y = 2x - 5 D. y = 5x - 2

OpenStudy (anonymous):

b

OpenStudy (anonymous):

What is the y-intercept of the line? y = -(1/2)x + 1 A. -1 B. -1/2 C. 1/2 D. 1

OpenStudy (anonymous):

a negative -1/2x plus 1 is 1 so D

OpenStudy (anonymous):

What is the y-intercept of the line? y = -3x A. -3 B. 0 C. 1 D. 3

OpenStudy (anonymous):

Thanks so much btw, you have no idea

OpenStudy (anonymous):

a

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

What is the y-intercept of the line? y = -3x – 4 A. -4 B. -3 C. 3 D. 4

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Write the equation of the line with the given slope, 1/2 and y-intercept, 7. A. y = (1/2)x - 7 B. y = (1/2)x + 7 C. y = 7x - ½ D. y = 7x + ½

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Write the equation of the line with the given slope, 6 and y-intercept, -8. A. y = -8x + 6 B. y = -6x + 8 C. y = 6x - 8 D. y = 8x - 6